OpenStudy (anonymous):

question on this integral, should i divide and what would it look like if i did

OpenStudy (anonymous):

\[\int\limits_{}^{}(3x^2+2x)/(x+1)\]

OpenStudy (amistre64):

3(-1)^2 +2(-1) ?=? 0

OpenStudy (amistre64):

factor the top; and split it up maybe?

OpenStudy (amistre64):

\[\frac{3x^2+2x}{x+1}=\frac{x(3x+2)}{x+1}=\frac{x}{x+1}....\] maybe not :)

OpenStudy (anonymous):

i did long dividion but i got left with a remainder or -x

OpenStudy (anonymous):

i am unsure of how i can proceed from there

OpenStudy (amistre64):

remainder should be 1 i thin

OpenStudy (anonymous):

well, -1 right, casue -x =-1

OpenStudy (amistre64):

integrate by parts; numerator as u and denominator as v?

OpenStudy (anonymous):

well, i guess i can do that, but i think i was suppose to incorporate susbtitution somehwere

OpenStudy (anonymous):

is there a different way to right the results after you divide poynomial?

OpenStudy (amistre64):

there is always a different way :) some gooder and some not so gooder

OpenStudy (anonymous):

cause i looked in a precalc book and it ws dividing (6x^2-26x+12)/ (x-4), and it wrote the results as : (6x-2)+((x)/(x-4)

OpenStudy (anonymous):

and if i could write it that way, then my life would be good

OpenStudy (anonymous):

cause i can just break it up

OpenStudy (anonymous):

thanks vicky for the link, but i cant look at that site, it will root my brain

OpenStudy (amistre64):

3x -1 + \(\cfrac{1}{x+1}\) ------------ x+1 | 3x^2 + 2x -3x^2 -3x ---------- -x +x+1 ------- 1

OpenStudy (amistre64):

lol .... that dint format like i wanted

OpenStudy (amistre64):

\[\int (3x -1 +\frac{1}{x+1})dx\]

OpenStudy (anonymous):

but that -1 should really be -x thought right

OpenStudy (amistre64):

no

OpenStudy (amistre64):

it would have looked better had the site parsed by carriage returns; but

OpenStudy (anonymous):